Determining of external funds needed.

n of External Funds Needed. Ina Corporation is thinking of purchasing a new his new machine, the company expects sales to increase from $8,000,000 to 3.3 Determinatio $10,000,000n thi new machine, the company expects sales to increase from $8,000,000 to any knows that its assets, accounts payable, and accrued expenses vary directly with mpanys profit margin on sales is 8 percent, and the company plans to pay 40 percent machine. With sales. T of its aft he co er-ta x earnings in dividends. The companys current balance sheet is given below. Balance Sheet 3,000,000 12,000,000 $15,000,000 $4,000,000 1,000,000 3,000,000 2,000,000 5,000,000 Total liabilities and net worth$15,000,000 Current assets Fixed assets Total assets Accounts payable Accrued expenses Long-term debt Common stock Retained earnings Prepare a pro forma balance sheet. Use the simple formula to determine the external funds needed by the company based on the answer in part (a) onditions 1) The profit margin rises from 8 percent to 10 percent. (a) (b) (o) Determine, using the simple formula, the external funds needed under each of the following The profit margin is 8 percent, but the dividend payout ratio is reduced from 40 percent to 20 percent. (d) Comment on the results from part (c).
Answer #1

All financials below are in $.

Sales last year, S0 = 8,000,000

Sales expected this year, S1 = 10,000,000

Growth rate in sales, g = S1 / S0 -1 = 10,000,000 / 8,000,000 -1 = 25%

Assets, accounts payables and accrued expenses are proportional to sales, hence they will grow by the same growth rate as Sales.

Part (a)

Please see the table below for pro forma statement.

Balance Sheet

Current

Pro forma

How proforma figure has been calculated

Current Assets

3,000,000

3,750,000

Current figure x (1 + g)

Fixed Assets

12,000,000

15,000,000

Current figure x (1 + g)

Total Assets

15,000,000

18,750,000

Sum of the above two items

Accounts Payables

4,000,000

5,000,000

Current figure x (1 + g)

Accrued expenses

1,000,000

1,250,000

Current figure x (1 + g)

Long term debt

3,000,000

5,020,000

Refer Note 2 below

Common stock

2,000,000

2,000,000

Same as current figure

Retained Earnings

5,000,000

5,480,000

See note 1 below

Total Liabilities & Net worth

15,000,000

18,750,000

Note 1:

Retained earnings in pro forma = Current retained earnings + change in retained earnings

Change in retained earnings = Expected Sales x Profit Margin x (1 - Dividend payout ratio) = 10,000,000 x 8% x (1 - 40%) = 480,000

Hence,  Retained earnings in pro forma = Current retained earnings + change in retained earnings = 5,000,000 + 480,000 = 5,480,000

Note 2:

Entire external funding is assumed to come through external debt. Hence, this is a plug figure after all other line items have been calculated.

Hence, pro forma long term debt = Total pro forma assets - Pro forma current liabilities - Pro forma accrued expenses - Proforma common stock - pro forma retained earnings = 18,750,000 - 5,000,000 - 1,250,000 2,000,000 - 5,480,000 = 5,020,000

Based on the pro forma statement, external fund needed = Incremental long term debt = Pro forma long term debt - Current long term debt = 5,020,000 - 3,000,000 = 2,020,000

Part (b)

Simple formula to determine the external funds needed

The standard formula for AFN is as shown below:

AFN = Additional Fund Needed

Ao So Lo So

Where, A0 = Total assets last year = $ 15,000,000

S0 = total sales last year = $ 8,000,000 mn

DeltaS = Change in sales = Expected sales - Last year's sales = 10,000,000 - 8,000,000 = 2,000,000

L0 = spontaneous liabilities = accounts payables + accruals = 4,000,000 + 1,000,000 = 5,000,000

PM = profit margin = 8%

S1 = Expected sales this year = 10,000,000

D = dividend payout ratio = 40%

Hence, AFN

15,000, 000 8,000,000 )× 2.000.000-( 5,000, 000 8,000, 000 ) × 2, 000, 000-0.08 × 10, 000. 000 (1- 0.40)

= $ 2,020,000

Part (c)

Part 1:

PM = profit margin = 8%

So, the AFN by the same formula as above but PM being 10% in stead of 8% this time

15,000, 000 8,000,000)× 2.000.000-( 5,000, 000 8,000, 000 ) × 2, 000, 000-0. 10 × 10, 000. 000 (1- 0.40)

= $ 1,900,000

Part 2:

D = 20%

So, the AFN by the same formula as above but D being 20% instead of 40% this time and PM being 8% as before,

15,000, 000 8,000,000 5,000, 000 )×2.000, 000-( ) 8,000, 000 × 2, 000, 000-0.08 × 10, 000. 000 1-0.20)

= $ 1,860,000

Part (d)

Internal accruals improve in both the cases. In part (1), the profit margins have improved from 8% to 10%, this has led to better internal accruals and hence, external funds needed have gone down. In part (2), internal accruals improved because payout reduced from 40% to 20%. This led to reduction on external funds needed.